OpenStudy (anonymous):

Estimate using compatible numbers. 9% of 72 a. 7 b. 14 c. 35 d. 70

OpenStudy (anonymous):

72*9/100 72*0.09= 6.48

OpenStudy (anonymous):

I don't understand.

OpenStudy (anonymous):

9% of 72 \[\frac{9}{100} \times 72 = 0.09 \times 72 = 6.48 \approx 7\]

OpenStudy (anonymous):

Ok. Thanks!!